OpenStudy (anonymous):

Solve please 2^(x) = 1/4

OpenStudy (anonymous):

\[2^{x}=0.25\]

OpenStudy (anonymous):

\[\log_{2} 2^{x}=\log_{2} 0.25\]

OpenStudy (anonymous):

\[x=\log_{2} 0.25\]

OpenStudy (anonymous):

What is the value?

OpenStudy (anonymous):

im not sure do you have a calculator with you? because i dont

OpenStudy (asnaseer):

you don't need a calculator for this

OpenStudy (anonymous):

oh right -2

OpenStudy (anonymous):

silly me

OpenStudy (asnaseer):

try and express 1/4 as a power of 2

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

but i am leading her through...?

OpenStudy (asnaseer):

@Lorena.isabel can you express 1/4 as a power of 2, i.e.:\[\frac{1}{4}=2^?\]

OpenStudy (anonymous):

ahhh....i see

OpenStudy (anonymous):

good moderator you are

OpenStudy (asnaseer):

thank you :)

OpenStudy (anonymous):

so the work above isn't enough?

OpenStudy (anonymous):

Whats the method?

OpenStudy (anonymous):

the very top

OpenStudy (anonymous):

lorena do you know what happens when you raise a number by a negative number?

OpenStudy (anonymous):

Oh i get it thanks

OpenStudy (asnaseer):

so what answer do you get for:\[\frac{1}{4}=2^?\]what should the ? be replaced by?

OpenStudy (anonymous):

-2

OpenStudy (asnaseer):

good, so now you know you can express:\[2^x=\frac{1}{4}\]as:\[2^x=2^{-2}\]and the solution should become apparent.

OpenStudy (anonymous):

Yeah thanks!

OpenStudy (asnaseer):

yw :)